Problem 13. A fair coin is tossed 10000 times. Find the probability that the number of tails is more or equal to 5050.
Answer #1

n = 10000

P(tail) = 0.5

\mu = n * p = 10000 * 0.5 = 5000

\sigma = sqrt(np(1 - p))

    = sqrt(10000 * 0.5 * 0.5)

    = 50

P(X > 5050)

= P(X > 5049.5)

= P((X - \mu )/\sigma> (5049.5 - \mu )/\sigma)

= P(Z > (5049.5 - 5000)/50)

= P(Z > 0.99)

= 1 - P(Z < 0.99)

= 1 - 0.8389

= 0.1611